### v2

# Why We Experience Society Differently: Intrinsic Dispositions as Drivers of Ideological Complexity in Adaptive Social Networks

Abstract
Understanding the emergence of inequality in complex systems requires attention to both structural dynamics and intrinsic heterogeneity. In the context of opinion dynamics, traditional models relied on static snapshots or assumed homogeneous agent behavior, overlooking how diverse cognitive dispositions shape belief evolution. While some recent models introduce behavioral heterogeneity, they typically focus on macro-level patterns, neglecting the unequal and individualized dynamics that unfold at the agent level. In this study, we analyze an adaptive social network model where each agent exhibits one of three behavioral tendencies—homophily, neophily (attention to novelty), or social conformity—and measure the complexity of individual opinion trajectories using normalized Lempel-Ziv (nLZ) complexity. We find that the resulting dynamics are often counterintuitive—homophilic agents, despite seeking similarity, become increasingly unpredictable; neophilic agents, despite pursuing novelty, stabilize; and conformic agents follow a U-shaped trajectory, transitioning from early stability to later unpredictability. More fundamentally, these patterns remain robust across diverse network settings, showing that internal behavioral dispositions — not external environment — primarily govern long-term opinion unpredictability. The broader implication is that individuals’ experiences of ideological volatility, uncertainty, or stability are not merely environmental, but endogenously self-structured through their own cognitive tendencies. These results establish a novel individual-level lens on opinion dynamics, where the behavioral identity of agents serves as a dynamical fingerprint in the evolution of belief systems, and gives rise to persistent disparities in dynamical experience within self-organizing social systems, even in structurally similar environments.
keywords:
adaptive social networks , behavioral heterogeneity , Lempel-Ziv complexity , homophily , social conformity , attention to novelty , self-organizing systems
journal: Physica A
[…]
[a]organization=Binghamton Center of Complex Systems, Binghamton University, State University of New York, addressline=, city=Binghamton, postcode=13902-6000, state=NY, country=USA
[…]
[b]organization=Waseda Innovation Lab, addressline=Waseda University, city=Tokyo, postcode=169-8050, country=Japan
1 Introduction
In today’s highly interconnected world, the increasing prevalence of extreme opinions and ideological polarization has become a defining characteristic of modern social systems [1, 2, 3, 4, 5, 6, 7]. While such polarization is commonly attributed to external triggers—such as political events or media manipulation—it can also emerge as a spontaneous outcome of social dynamics driven by human-centric factors [8, 9]. The advent of modern information communication technology has further intensified these dynamics by enabling preferential selection of information sources, often leading to the formation of “social bubbles” and echo chambers [10, 11, 12]. As societies grapple with issues such as globalization, immigration, and cultural diversity, individuals and groups may retreat into more homogeneous communities as a means of preserving their sense of identity and security.
To understand the formation of extreme opinions and the structural evolution of fragmented societies, researchers have studied phase transitions between connected and fragmented states in adaptive social networks [13, 14, 15, 16]. Within these adaptive social networks, individuals adjust their social ties and opinions based on local interactions, producing complex, self-organizing patterns of social fragmentation and ideological entrenchment. Traditional bounded confidence models have demonstrated that polarization can emerge even under moderate openness to differing views [17, 18]. Prior research has shown that heterogeneity in update policies can significantly influence the formation of extremist communities and network connectivity [19, 20].
A key limitation of much of the existing literature on opinion dynamics lies in its reliance on static snapshots of network states or opinion distributions. These approaches—often based on modularity, clustering, or structural detection of polarization—offer only a momentary view of ideological divisions, failing to account for how opinions and network structures evolve over time. In response, dynamic models of social networks were developed to capture temporal evolution; however, these models typically assumed homogeneous behavioral rules across agents, neglecting the diversity of cognitive dispositions that shape real-world social behavior [10, 21].
To address this, more recent studies introduced behavioral heterogeneity into dynamic opinion models [9, 22, 23]. While this marked an important advancement, such analyses have largely focused on aggregate or system-level outcomes, such as overall polarization or network connectivity. These macro-scale metrics, while informative, obscure the potentially rich and unequal dynamical behaviors at the level of individual agents.
This study addresses that gap by explicitly analyzing individual-level opinion dynamics within a behaviorally heterogeneous adaptive social network. By tracking the temporal unpredictability of each node’s opinion trajectory, we uncover how internal behavioral dispositions—not just external network structure—govern the complexity of ideological evolution. This focus on individual dynamical fingerprints represents a novel approach to understanding polarization and social fragmentation in adaptive systems.
We employ the normalized Lempel-Ziv (nLZ) complexity to quantify the variability of node opinion trajectories in an adaptive social network model [24, 25]. We study the adaptive social network model developed in [8] that explains how extreme ideas can spontaneously arise based on the interplay between three factors: homophily, attention to novelty (neophily), and social conformity. Here, it was found that, while homophily reinforces segregation by strengthening ties between like-minded individuals, neophily encourages diversity by fostering interactions with those who hold novel viewpoints. Social conformity further complicates these dynamics, as it can either promote ideological convergence or drive extremization depending on network conditions. nLZ complexity, here, can be understood as the characterization of the unpredictability in the opinion-evolution trajectory of nodes in this adaptive social network. A node undergoing a more rapid, diverse, and extreme set of opinions can be said to have an experience of a more dynamic and unpredictable nature as compared to that of a node undergoing a more constrained and redundant set of opinions in the network. The use of this measure, therefore, allows us to understand the relationship between the inherent unpredictability in opinion trajectories of nodes in these adaptive networks with respect to their internal dispositions (homophily, neophily, and social conformity).
By analyzing nLZ trends across three experimental scenarios (described in the following section), our objective is to identify faction-specific complexity patterns of opinion trajectories and examine how node-level behavioral tendencies shape the dynamical complexity of opinion evolution.
2 Methods
2.1 The model
The original model [8] describes an adaptive social network with n nodes, with each node i having its own opinion $x_{i}\in\mathbb{R}$. The edge weight $w_{ij}$ denotes flow of information from source node j to target node i. The dynamics of the network, as indicated by evolving nodal opinions and edge weights are described by the following set of equations:
$\dfrac{dx_{i}}{dt}$ $=c(\langle x\rangle_{i}-x_{i})+\epsilon$ (1)
$\dfrac{dw_{ij}}{dt}$ $=hF_{h}(x_{i},x_{j})+aF_{a}(\langle x\rangle_{i},x_{j})$ (2)
$\langle x\rangle_{i}$ $=\dfrac{\sum_{j}w_{ij}x_{j}}{\sum_{j}w_{ij}}$ (3)
Here, $\langle x\rangle_{i}$ is the weighted average of opinions in node i’s neighborhood, i.e., the social norm as perceived by node i. c and $\epsilon$ are social conformity and noise terms that impact the evolution of node i’s opinion. In the second equation, h and a are the homophily and attention to novelty factors that impact the change in the edge weight for the edge directed from node j to node i, in other words, the degree of closeness between the two nodes i and j. Here, $F_{h}$ and $F_{a}$ are behavioral functions that determine the degree of change in edge weights. These functions are adopted from the original study and are described below:
$F_{h}(x_{i},x_{j})$ $=\theta_{h}-|x_{i}-x_{j}|$ (4)
$F_{a}(\langle x\rangle_{i},x_{j})$ $=|\langle x\rangle_{i}-x_{j}|-\theta_{a}$ (5)
$\theta_{h}$ and $\theta_{a}$ are constants that define the function values when the arguments are equal. Essentially, $F_{h}$ designates the strengthening of edge connecting j to i when nodes i and j have similar opinions. $F_{a}$ designates the strengthening of the same edge when node j has an opinion different from that of the perceived social norm of node i (i.e., novelty). In these simulations, $w_{ij}$ is bound to be non-negative and the edge weights are set to zero when they go negative.
Figure 1: Functional shapes for the behavioral functions $F_{h}$ (left) and $F_{a}$ (right) for $\theta_{h}=\theta_{a}=0.3$ (adopted from [8]). The cyan planes indicate zero level.
2.2 Lempel-Ziv Complexity
Lempel-Ziv (LZ) complexity measures a sequence’s predictability by assessing its compressibility [24]. It identifies repeated substrings within the sequence; the more unique substrings found, the greater the unpredictability. High LZ sequences are harder to compress as they contain fewer repeating patterns, while low LZ sequences compress more readily due to greater redundancy. Therefore, the efficiency of compression reflects the sequence’s richness: more compressible sequences signify greater predictability, while resistance to compression indicates greater unpredictability.
To calculate the LZ complexity, we traverse a string from left to right, adding each new substring to a dictionary. If a substring repeats, we extend the search span to find a new unique substring. For instance, in “01100101101100100110”, we encounter “0” and “1,” each added to the dictionary. Then, upon encountering another “1” we extend the search span to two, to “10”, which is added. Continuing in this manner, the string is parsed as 0•1•10•010•1101•100100•110 [26]. Note that the final substring may not be unique. The LZ complexity of this sequence is equal to the dictionary’s length, representing the number of unique substrings. In this case, it is 7.
In this study, we use the normalized Lempel-Ziv (nLZ) complexity which, essentially, makes the measure independent of the length of the sequence (since LZ is clearly correlated with sequence length) [25]. This is done as:
$\text{{nLZ}}=\dfrac{LZ}{\left(\dfrac{\text{{n}}}{\text{log {n}}}\right)}$ (6)
where n is the length of the sequence.
This is justified, since it has been shown that the LZ complexity for a sequence of uniformly distributed characters approaches nlog $\dfrac{\text{{n}}}{\text{log {n}}}$ as n becomes arbitrarily large [24].
For our model, we compute the nLZ complexity for each node’s time series of opinions (until a given time step) and average this across all nodes of a given type (depending on the update policy - highly homophilic/neophilic/conformic) to obtain a measure of the predictability of opinion trajectory for nodes of that faction.
2.3 Discretization of the model
There are two types of discretization that are implemented in our model, temporal and opinion.
Temporal: The ordinary differential equation (ODE) based model is discretized using the Euler-forward approach. More specifically, the states for each node are updated in discrete time steps of size 0.1. Let t be the time step being computed. $t=0$ implies a simulation time of 0; $t=10$ implies a simulation time of 1; $t=50$ implies a simulation time of 5, and so on.
Opinion: The nLZ complexity can only be computed for a discrete sequence of characters. In our ODE model, the opinion states of nodes are continuous. The opinion states are, therefore, discretized by implementing a mean-standard deviation binning approach. Numerical simulations of the model for a number of initial conditions and parameter settings revealed the following:
  • 1.
    For networks with parameter settings biased towards fragmentation (high $h$), the mean standard deviation of states ($\sigma$) tended to be greater than or close to 2.
  • 2.
    For networks with parameter settings biased towards strong connectivity (high $a$ $c$), the mean $\sigma$ tended to be less than or close to 1.
  • 3.
    For networks with a uniform distribution of parameter settings, the mean $\sigma$ of states tended to be close to 1.5.
[…]
Note that results remained consistent up to bin sizes of 1. Beyond this, the effect of larger bins yielded inconsistent values of nLZ complexity.
2.4 Experimental setup
Three scenarios were devised, each an extension of the previous, to explore the space of possibilities in the adaptive social network model. These are described below.
  • 1.
    Scenario 1: Create “pure” networks comprising of agents, all biased toward one of the update policies.
    • (a)
      Homophilic networks: a, c $\sim$ $N(0.05,0.025)$; h $\sim$ $N(0.25,0.025)$
    • (b)
      Neophilic networks: h, c $\sim$ $N(0.05,0.025)$; a $\sim$ $N(0.25,0.025)$
    • (c)
      Conformic networks: h, a $\sim$ $N(0.05,0.025)$; c $\sim$ $N(0.25,0.025)$
  • 2.
    Scenario 2: Generate 50-50 “dual faction” networks comprising of two equally proportioned factions each biased toward one of the three update policies.
    • (a)
      This creates three types of networks where we observe network evolution as a result of interactions between two competing factions in the population (h vs a, a vs c, c vs h).
    • (b)
      Node parameterization is based on the same distributions as Scenario 1 (depending on the faction).
  • 3.
    Scenario 3: Create networks of mixed agents (with uniformly distributed node parameters).
    • (a)
      a, c, h $\sim$ $U(0.05,0.03)$
In Scenario 1, pure networks are created where all agents share the same bias—resulting in homophilic, neophilic, or conformic networks. In Scenario 2, dual faction networks are formed with two competing biases (h vs a, a vs c, c vs h) to study how interactions between two factions affect opinion complexity trends. Finally, Scenario 3 uses a mixed setup with uniformly distributed biases, providing a fully heterogeneous environment. This approach systematically examines how opinion complexity trends vary from homogeneous settings (where all agents are of the same type) to heterogeneous settings (comprising of a mixed pool of different types of agents).
All parameter distributions have been adopted from the original study (exploring the impact of heterogeneity in update policies) to ensure consistency [20]. Each network consists of 300 nodes. 10 networks each were generated for the three scenarios (10 each for the three pure networks of Scenario 1; 10 each for the three dual faction networks of Scenario 2; 10 mixed networks for Scenario 3), and simulated for 3000 time-steps.
3 Results
Figures 23, and 4 illustrate the evolution of nLZ complexity over time across all three experimental scenarios. nLZ complexity is computed for each node’s opinion trajectory over incrementally increasing time windows (from $t=0$ to $t=3000$ in steps of 300), and results are averaged within each node type to assess characteristic patterns.
Figure 2: nLZ vs t for the three pure networks in Scenario 1. Homophilic (left) vs Neophilic (center) vs Conformic (right) node nLZ complexities (with 95% confidence bounds) are shown in distinct colors. The nLZ value is computed over increasing time intervals to capture changes in opinion trajectorial unpredictability over time.
Figure 3: nLZ vs t for the three 50-50 networks in Scenario 2 (with 95% confidence bounds). Homophily vs Neophily (left); Neophily vs Conformity (center); Conformity vs Homophily (right).
Figure 4: nLZ vs t for the mixed networks in Scenario 3 (with 95% confidence bounds).
3.1 Distinctive complexity profiles by node type
The analysis reveals robust, reproducible, and behavior-specific complexity signatures for homophilic, neophilic, and conformic nodes across all scenarios. Homophilic nodes consistently exhibit a steadily increasing nLZ trend, indicating progressively more unpredictable opinion dynamics. While this appears counterintuitive—given that homophilic individuals are presumed to seek stability through similarity—the model reveals that their persistent reinforcement of like-mindedness may paradoxically lead to long-term instability. Local fluctuations, minor disagreements, and evolving neighborhood structures introduce compounding variability, thereby driving up the unpredictability of their trajectories.
In contrast, neophilic nodes maintain a low and stable nLZ complexity profile over time, across all network settings. Although these nodes actively seek novelty, their trajectories stabilize early in the simulation, indicating that exposure to diverse opinions induces a saturation in unpredictability. The system essentially “absorbs” complexity early on, allowing neophilic nodes to maintain consistent behavior despite the dynamic social environment.
Conformic nodes display a U-shaped nLZ trend. Initially, conformity suppresses diversity, resulting in predictable and redundant opinion dynamics. However, as the network evolves, subtle divergences and local inconsistencies emerge, leading to renewed unpredictability and a rise in nLZ complexity over time. This dual-phase behavior suggests that social conformity can foster both short-term homogeneity and long-term instability, depending on the stage of network evolution.
3.2 Internal behavioral rules as drivers of complexity
A striking and novel observation in this study is the robustness and consistency of these nLZ complexity trends across all three experimental scenarios—pure, dual-faction, and fully heterogeneous networks. Regardless of the surrounding social environment, each node type maintains a unique and reproducible complexity signature. This indicates that the internal behavioral predispositions of nodes—homophily, neophily, or conformity—are the dominant determinants of the complexity of their opinion dynamics, rather than the external network composition or factional mixture.
This finding has important implications. It suggests that the perceived complexity of an individual’s social experience is primarily a reflection of their own internal tendencies, rather than the social environment in which they are embedded. In other words, individual-level experiences of opinion instability and unpredictability are less about external stimuli and more about internal behavioral inclinations. For example, a homophilic individual may experience greater volatility in their worldview not due to a chaotic society, but because their own persistent quest for similarity engenders systemic instability.
This insight generalizes to broader psychological and sociological interpretations: when individuals experience a turbulent or unsatisfactory social existence, they may benefit more from reflecting on their internal cognitive dispositions than attributing causality solely to external conditions. The adaptive social network model thus captures a deeper truth—that one’s subjective social complexity may be intrinsically rooted in temperament, not circumstance.
3.3 Counterintuitive behavioral dynamics
The behavioral patterns observed in the model are frequently unintuitive and challenge common assumptions about the psychological correlates of homophily, novelty-seeking, and conformity:
  • 1.
    Homophilic nodes, despite a drive for consistency and alignment, demonstrate increasingly unpredictable opinion trajectories. This paradox highlights the destabilizing long-term consequences of excessive self-reinforcement within homogeneous subgroups.
  • 2.
    Neophilic nodes, expected to be chaotic due to constant novelty-seeking, display stable and plateauing complexity. Early saturation from diverse exposure appears to buffer further volatility.
  • 3.
    Conformic nodes exhibit biphasic dynamics, with initial predictability followed by delayed unpredictability. Early convergence is gradually undone by fragmentation and local divergences, underscoring the temporal fragility of social consensus.
3.4 Interaction effects in heterogeneous networks
In fully heterogeneous networks (Scenario 3), where all three behavioral types coexist, interaction effects are evident, though subtle. Homophilic nodes still show the highest and increasing nLZ trend, but with a slightly reduced slope compared to pure or dual-faction settings. Neophilic nodes maintain a plateau but exhibit a minor downward shift in their baseline complexity, while conformic nodes continue to follow a U-shaped curve that is now elevated in magnitude and less pronounced in curvature.
Notably, conformic nodes eventually surpass neophilic nodes in unpredictability—a reversal of trends seen in less diverse settings. These shifts reflect the moderating or amplifying effects of inter-faction interactions. In mixed environments, exposure to a broader range of behavioral influences tempers extreme tendencies and redistributes unpredictability across node types. Yet even under these interaction effects, the core identity of each node type remains intact, reinforcing the finding that internal behavioral parameters are primary drivers of long-term complexity outcomes.
4 Conclusions
This study investigated the dynamics of opinion evolution in adaptive social networks by leveraging normalized Lempel-Ziv (nLZ) complexity as a measure of opinion trajectory unpredictability. Through simulations across three experimental settings—pure, dual faction, and heterogeneous—we identified robust, behavior-specific opinion complexity signatures associated with nodes biased toward homophily, neophily, and social conformity. Notably, these signatures were often counterintuitive: homophilic nodes—despite seeking similarity—became increasingly unpredictable; neophilic nodes—despite courting novelty—stabilized early; and conformic nodes followed a U-shaped progression from initial stability to later unpredictability.
A central and novel contribution of this work is the discovery that the unpredictability of an individual’s opinion trajectory is remarkably consistent across varying network compositions, and is instead primarily governed by internal behavioral predispositions. Whether situated in homogeneous or heterogeneous environments, each node type maintained its characteristic nLZ complexity profile. This suggests that intrinsic behavioral factors—such as a preference for similarity, novelty, or conformity—play a dominant role in shaping one’s social experience, often overriding the influence of external social structure.
This insight has broader implications beyond the technical findings. It implies that the complexity of one’s ideological or social trajectory is not merely a reflection of environmental volatility, but rather an emergent property of individual temperament. For example, a homophilic individual may experience growing instability not because of external disruption, but because their behavioral orientation fosters feedback loops that amplify unpredictability. In contrast, a neophilic agent might quickly encounter and assimilate novelty, leading to stabilized opinion dynamics. These patterns reveal that the perceived chaos or cohesion of one’s social experience may be largely self-structured.
From a societal perspective, this finding challenges common assumptions that attribute polarization and ideological volatility primarily to structural or environmental causes. If unpredictability in opinion dynamics arises chiefly from within, then interventions aimed at promoting social stability or cohesion may need to address underlying cognitive and behavioral dispositions, rather than focusing solely on reforming external conditions. This dual approach—targeting both structure and behavior—could yield more effective strategies for mitigating polarization and fragmentation.
Our study also highlights the persistence of individual behavioral “fingerprints” even under strong interaction effects. While heterogeneous environments do induce mild shifts in complexity trends, they do not erase the fundamental patterns tied to each behavioral type. This underscores the resilience and primacy of internal behavioral rules in shaping long-term opinion dynamics.
Finally, the uniqueness of our approach lies not merely in using nLZ complexity, but in focusing on temporal shifts and individualized complexity trajectories in behaviorally diverse networks. Future work could extend this perspective by incorporating complementary complexity measures—such as entropy, fractal dimension, or recurrence quantification—to uncover additional facets of opinion dynamics. More broadly, these results suggest a need for a paradigm shift in opinion dynamics research—one that centers individual behavioral identity as a foundational element in the emergence of social complexity.